# High-Fidelity Scientific Simulation Surrogates via Adaptive Implicit Neural Representations

### v1

@@ 1 Introduction
Accurately simulating complex physical systems is vital for scientific discovery, but high-fidelity simulations are often prohibitively expensive to run at scale [6, 38, 49]. Surrogate models offer a practical alternative by learning from simulation outputs or observational data to deliver fast and accurate predictions [4, 5, 8, 17, 24, 39, 41, 42]. This enables researchers to efficiently explore scientific phenomena, test hypotheses, and support decision-making. For example, climate scientists can use surrogate models to study how ocean temperatures respond to varying wind stress without repeatedly running costly simulations.
@@ 3.2 Challenges
Explicit representations for INRs. Recent approaches further improve the representational power of INRs by incorporating explicit feature encodings [4, 14, 15, 44, 45] into the MLP-based framework, as illustrated in Figure 2. These encodings often employ structured representations, such as feature grids, which store learnable feature vectors at discrete locations (vertices). For example, for three-dimensional input coordinates, a feature grid $\mathcal{F}\in\mathbb{R}^{{N_{1}}\times{N_{2}}\times{N_{3}}\times{D}}$ can be utilized. Here, ($N_{1}$,$N_{2}$, $N_{3}$) defines the grid resolution, and $D$ is the dimension of the feature vectors stored at each vertex. To obtain the feature vector $z^{(x)}$ of an arbitrary input coordinate $x\in\mathbb{R}^{d}$, the coordinate $x$ is first mapped from its original spatial domain to a normalized feature space $[-1,1]^{3}$ through a linear projection $\phi$. Then, $z^{(x)}$ is retrieved by interpolating the features stored at the neighboring vertices of $x$ within the grid $\mathcal{F}$. The feature interpolation process can be defined as:
@@ 4.1 Adaptive Encodings via Cross-Attention
Specifically, to effectively incorporate the simulation parameters into the feature retrieval process, each parameter $s\in S$ is first individually embedded into a feature vector ${z_{s}}^{(p_{j})}$ using one-dimensional feature vectors. These embeddings are then combined through an element-wise (Hadamard) product [8, 14] to better preserve the signal from each of the simulation parameters. As shown in Figure 3, the output parameter-conditioned embedding is then concatenated with each retrieved value in the memory bank and further processed by the adapter $f_{\theta_{\text{A}}}$:
@@
This residual connection here aims to better preserve the original spatial information while explicitly adapting the features for the given simulation parameters $p_{j}$.
@@ Appendix D Qualitative Results
In the main paper, we demonstrated the rendered images of the Nyx test dataset. In this section, we provide additional qualitative results for the MPAS-Ocean and CloverLeaf3D datasets. All images were rendered using ParaView, an open-source platform built on the Visualization Toolkit (VTK) libraries.
@@ Appendix F Alternative Approach for Expert Aggregation
The proposed FA-INR employs a Mixture of Experts (MoE) architecture, specifically utilizing Top-2 expert selection, to effectively balance training efficiency with prediction accuracy. This section discusses an alternative approach to arrange the experts.
Instead of making the Top-2 expert selection, this alternative method forwards each input to all expert encoders simultaneously. The outputs from every expert are then concatenated into a single large feature vector. Finally, this aggregated vector is projected back to the original feature dimension before being processed by the feature decoder. As presented in Table 6, the original design of MoE with Top-2 expert selection outperforms the concatenation strategy, particularly achieving a 34% improvement in the MD metric. We hypothesize that the performance degrades for the concatenation-based strategy because, in contrast to the MoE framework, where experts can develop specialized knowledge, the concatenation method forces every expert to process all input data. This approach may lead to duplication in the learned representations across experts, substantially increasing the training time and further limiting the prediction accuracy.
Metric MoE Concatenation
#Experts 10 10
@@
Table 6: Comparison of encoder expert utilization strategies (i.e., Top-2 selection vs. all-expert concatenation) on the MPAS-Ocean dataset.
@@ Appendix G Generalization in Spatial Domain
While our proposed FA-INR surrogate primarily focuses on improving generalization across unseen simulation parameters (i.e., as demonstrated in subsection 5.2), we further evaluate its ability to generalize to unseen spatial locations. Specifically, during training, we randomly select 70% of spatial coordinates within each of the 70 ensemble members, reserving the remaining 30% as a test set.
The quantitative results across all baselines are summarized in Table 7. Overall, FA-INR consistently outperforms baseline INR models, achieving an average improvement of 13.67% in PSNR and 42.59% in MD. Interestingly, Explorable-INR exhibits relatively poor spatial generalization compared to its performance in the main experiments. In contrast, K-Planes achieves competitive performance on both metrics, indicating that although K-Planes shows difficulty in generalization across simulation parameters, its multi-resolution feature planes are particularly effective in capturing spatial information, despite having significantly more model parameters than other methods.
@@
Table 7: Comparing the generalization abilities in the spatial domain on Mpas-Ocean dataset.

### v2

@@ 1 Introduction
Accurately simulating complex physical systems is vital for scientific discovery, but high-fidelity simulations are often prohibitively expensive to run at scale [7, 39, 50]. Surrogate models offer a practical alternative by learning from simulation outputs or observational data to deliver fast and accurate predictions [6, 9, 18, 25, 40, 42, 43]. This enables researchers to efficiently explore scientific phenomena, test hypotheses, and support decision-making. For example, climate scientists can use surrogate models to study how ocean temperatures respond to varying wind stress without repeatedly running costly simulations.
@@ 3.2 Challenges
Explicit representations for INRs. Recent approaches further improve the representational power of INRs by incorporating explicit feature encodings [5, 15, 16, 45, 46] into the MLP-based framework, as illustrated in Figure 2. These encodings often employ structured representations, such as feature grids, which store learnable feature vectors at discrete locations (vertices). For example, for three-dimensional input coordinates, a feature grid $\mathcal{F}\in\mathbb{R}^{{N_{1}}\times{N_{2}}\times{N_{3}}\times{D}}$ can be utilized. Here, ($N_{1}$$N_{2}$, $N_{3}$) defines the grid resolution, and $D$ is the dimension of the feature vectors stored at each vertex. To obtain the feature vector $z^{(x)}$ of an arbitrary input coordinate $x\in\mathbb{R}^{d}$, the coordinate $x$ is first mapped from its original spatial domain to a normalized feature space $1,1]^{3}$ through a linear projection $\phi$. Then, $z^{(x)}$ is retrieved by interpolating the features stored at the neighboring vertices of $x$ within the grid $\mathcal{F}$. The feature interpolation process can be defined as:
@@ 4.1 Adaptive Encodings via Cross-Attention
Specifically, to effectively incorporate the simulation parameters into the feature retrieval process, each simulation condition is represented by a vector $p_{j}$, where each element $s\in S$ denotes an individual scalar parameter within $p_{j}$. Each scalar $s\in S$ is first individually embedded into a one-dimensional feature vector ${z_{s}}^{(p_{j})}$. These scalar-level embeddings are then combined via an element-wise (Hadamard) product [9, 15] to form a unified parameter embedding ${z}^{(p_{j})}$, which better preserves the signal from each individual parameter. As shown in Figure 3, this integrated embedding is then concatenated with each retrieved value in the memory bank and further processed by the adapter $f_{\theta_{\text{A}}}$:
@@
This residual connection is designed to better preserve the original spatial information while explicitly adapting the features based on the specific simulation parameters $p_{j}$.
@@ Appendix D Qualitative Results
In the main paper, we demonstrated the rendered images of the Nyx test dataset. In this section, we provide additional qualitative results for the MPAS-Ocean and CloverLeaf3D datasets. All images were rendered using ParaView [2], an open-source platform built on the Visualization Toolkit (VTK) libraries.
@@ Appendix F Alternative Approach for Expert Aggregation
The proposed FA-INR utilizes a Mixture of Experts (MoE) architecture with Top-2 routing strategy, effectively balancing training efficiency and prediction accuracy. This section discusses two alternative expert arrangement strategies.
First, instead of making the Top-2 selection, one alternative forwards each input to all expert encoders simultaneously. The outputs from all experts are concatenated into a single large feature vector, which is then projected back to the original feature dimension before being processed by the feature decoder. As presented in Table 6, the original MoE design with Top-2 expert selection outperforms this concatenation strategy, particularly achieving a 34% improvement in the MD metric. We hypothesize that the performance degradation arises because, unlike the MoE framework where experts can develop specialized knowledge, the concatenation method forces every expert to process all inputs. This can lead to duplicated representations across experts, substantially increasing training time while limiting prediction accuracy.
Second, we evaluated a weighted sum strategy over all experts, where the outputs from all experts are aggregated using learned weights instead of sparse gating. Although this approach provided a slight improvement in overall accuracy compared to Top-2 routing, as shown in Table 6, it substantially increased inference time. Specifically, on a single A100 GPU, the Top-2 routing strategy required only $1.70$ seconds per test field, whereas the weighted sum strategy took $7.42$ seconds. These results highlight that activating more experts per spatial input significantly increases computational overhead. Therefore, our sparsely-gated Top-2 routing approach offers a more practical balance between computational efficiency and model accuracy.
Metric MoE Concatenation Weighted Sum
#Experts 10 10 10
@@
Table 6: Comparison of encoder expert utilization strategies (i.e., Top-2 selection vs. all-expert concatenation vs. weighted sum over all experts) on the MPAS-Ocean dataset.
@@ Appendix G Generalization in Spatial Domain
While our proposed FA-INR surrogate is primarily designed to improve generalization across unseen simulation parameters (as demonstrated in subsection 5.2), we further evaluate its ability to generalize to unseen spatial locations. This evaluation provides complementary insight under a different generalization setting.
Specifically, during training, we randomly select 70% of the spatial coordinates within each of the 70 ensemble members and reserve the remaining 30% as a test set. This setting keeps the simulation parameters fixed and only evaluates generalization over spatial locations.
The quantitative results across all baselines are summarized in Table 7. Overall, FA-INR consistently outperforms baseline INR models, achieving an average improvement of 13.67% in PSNR and 42.59% in MD. Interestingly, Explorable-INR exhibits relatively poor spatial generalization compared to its performance in the main experiments. In contrast, K-Planes achieves competitive performance on both metrics, indicating that although K-Planes shows difficulty in generalization across simulation parameters, its multi-resolution feature planes are particularly effective for modeling spatial variations. However, this comes at the cost of significantly more model parameters.
@@
Table 7: Evaluation of generalization to unseen spatial locations on the MPAS-Ocean dataset. Note that this spatial generalization task is distinct from the task in our main paper, which focuses on generalization across unseen simulation parameters.